A monopolist faces the demand curve Q=60-P/2. The cost function is C=Q2 . Find the output that maximizes this monopolist's profit. What are the prices at point and that output? Find the elasticity of demand at the profit maximizing output.


1
Expert's answer
2021-03-11T08:02:33-0500


Q = 60 - P/2 or P = 120 - 2Q.

Monopolist profit-maximizing quantity is produced, when MR = MC.

"MR = TR' = (P*Q)' = (120Q - 2Q^{2})' = 120 - 4Q,"

"MC = C' = (Q^{2})' = 2Q."

120 - 4Q = 2Q,

Q = 20 units.

P = 120 - 2*20 = 80.

"Ed = -b*P\/Q = -0.5*80\/20 = -2,"

so the demand is elastic.